OpenStudy (anonymous):

what is the area of a rectangle that is 4x4 1/2 (please explain your answer) what is the area of a rectangle that is 3 1/2x6 1/2? (please explain your answer)

OpenStudy (pinklion23):

is this part two of your assessment?

OpenStudy (pinklion23):



OpenStudy (pinklion23):

are you there?

OpenStudy (pinklion23):

I'm not giving you the answer because I'm not aloud to

OpenStudy (anonymous):

it doesn't show a image.....

OpenStudy (pinklion23):

ok well to find area you have to do lw=a so 4*4.5=A you understand

OpenStudy (anonymous):

yea n that's what i did yesterday and i got both of them wrong....

OpenStudy (pinklion23):

humf well to find area that is what you have to do so 4*4.5=18 is that what you got for 1

OpenStudy (anonymous):

no what i did was turned 4 1/2 into a whole number and just multiplied the whole numbers....

OpenStudy (pinklion23):

well it does say round to the nearest whole now does it there for you must keep the numbers as they are all you must do is convert 1/2 to 0.5 then you get 4*4.5=18

OpenStudy (anonymous):

ok thanks

OpenStudy (pinklion23):

no problem so do you think you can do number 2 by your self?

OpenStudy (anonymous):

idek.....

OpenStudy (pinklion23):

ok well we can work together to get the right answer

OpenStudy (pinklion23):

what is the area of a rectangle that is 3 1/2x6 1/2? (please explain your answer)

OpenStudy (pinklion23):

so what do you need to do first?

OpenStudy (pinklion23):

you there

OpenStudy (pinklion23):

may i have medal?